OpenStudy (anonymous):

What is a simpler form of the following expressions? (7p – 8 )(7p + 8 )

OpenStudy (anonymous):

Difference of squares. \[(a-b)(a+b) = a^2-b^2\]In this case \(a=7p\) and \(b= 8\). So we get \[(7p)^2-8^2\]Which comes out to be:\[49p^2-64\]When simplified further.

OpenStudy (anonymous):

You could also use foil, but I think they want you to recognize the difference of squares.

OpenStudy (anonymous):

7p+-8